I have to combine like terms or equation. I want to know did I do this promblem right 4+4+9x+10x=8+19x
Answer by nyc_function(2645)

Simply combine constants with constants and terms with terms.
4+4+9x+10x=8+19x
8 + 19x = 8 + 19x
Do you see what's going on here?
There is no solution to this linear equation.
When we combine like terms at this point, we get 0 = 0, which means NO SOLUTION.